A poll is given, showing 80% or in favor of a new building project.


If seven people are chosen at random, what is the probability that exactly 6 of them favor the new building project.
Answer #1

answer=0.3670

here we use binomial distribution with parameter n=7 and p=0.8

for Binomial distribution ,P(X=r)=nCrpr(1-p)n-r  

probability that exactly 6 of them favor=P(X=6)=7C6(0.8)6(1-0.8)7-6=7*0.2621*0.20=0.3670

we can also use ms-excel command =BINOMDIST(6,7,0.8,0) for the above answer)